How do I solve this IVP by using the method of Laplace Transforms?

Preferably anyone who knows Differential Equations very well.

0 if 0 st <3 y + 2y + 5y = g(t) where g(t) = {t-3 if 3 st<8 0 if t28 y(0) = -1, y(0) = 1
Answer #1

Pago Ma: Data: given ++ 457 - 94) ost 23 f £78 wesolve for so yl +2y + 5y = 0 opply tapluce transforum bothside LCyll] TalNow get) a t-3 four M tay & 5y = 4-3 [CY!!+22(40)*+ 5 LC4) = 164-3) 403) (${f2545) +(5+1) s 3 S² S 715) + (StD stas +5 (6²)